write the equation of the line that passes through (-3,-2) and (1,-2) in the form y=mx+b
Answer by mananth(16949)

x1 y1 x2 y2
-3 -2 1 -2

slope m = (y2-y1)/(x2-x1)
( -2 - -2 )/( 1 - -3 )
( 0 / 4 )
m= 0
The required line passes through both the points. So we select one point and plug it with the slope
Plug value of the slope and point ( -3 , -2 ) in
Y = m x + b OK
-2.00 = 0 * -3 + b
-2.00 = -0 + b
b= -2 + -2
b= 0
So the equation will be
Y = -2
m.ananth@hotmail.ca